# IPFormer: Visual 3D Panoptic Scene Completion with Context-Adaptive Instance Proposals

### v2

@@ 1 Introduction
Figure 1: Comparison of query initialization for Panoptic Scene Completion (PSC). Existing methods, e.g. Symphonies symphonies , randomly initialize instance queries and incorporate context-awareness during training. However, these queries retain their static nature at test time, as they are shared across all inputs. Our method IPFormer initializes them as instance proposals, which preserve context-adaptivity at test time, effectively aggregating directed features for improved PSC performance. Due to the the novelty of vision-based PSC and the absence of established baselines, we apply DBSCAN ester1996_dbscan clustering to Symphonies’ SSC output to retrieve its individual instances.
@@
Panoptic Scene Completion provides a holistic scene understanding that can serve applications like autonomous driving and robotics by reconstructing volumetric geometry from sparse sensor data and assigning meaning to objects in the scene. Such holistic scene understanding requires both accurate geometry and semantic information about a scene. Historically, these tasks have been treated separately and evolved independently due to the distinct nature of geometric reconstruction and semantic interpretation, each requiring specialized algorithms and data representations. Despite the inherent interconnection between geometry and semantics, their separation hindered unified scene understanding. Classical structure-from-motion evolved into simultaneous localization and mapping (SLAM) systems for accurate reconstructions mur2015orb ; engel2014lsd ; koestler2022tandem , while deep learning facilitated monocular depth estimation (MDE) methods godard2017unsupervised ; godard2019digging ; wimbauer2021monorec , which are restricted to visible surfaces. In contrast, Scene Completion refers to the process of inferring the complete 3D structure of a scene, including both visible and occluded regions, from partial observations wimbauer2023behind ; han2024boosting .
@@
Apart from geometry, semantic image segmentation divides an image into semantically meaningful regions and labels them yu2015multi ; chen2017deeplab . Panoptic segmentation kirillov2019panoptic ; kirillov2019panoptic2 ; porzi2019seamless extends this approach by segmenting instances hariharan2014simultaneous for things (countable objects like person or car) while maintaining semantic segmentation for stuff (amorphous regions like road or vegetation). This unified approach is vital for comprehensive scene understanding, as it distinguishes individual object instances while preserving semantic context for background elements, enabling applications like autonomous navigation. The recent trend in scene understanding has been to unify these tasks into Semantic Scene Completion (SSC) hayler2024s4c and Panoptic Scene Completion (PSC) pasco . SSC integrates geometric completion with semantic labeling, predicting both the 3D structure and semantic categories, represented as a 3D grid composed of voxels. PSC builds on this by adding instance-level segmentation for each voxel, thus distinguishing between individual objects. While PSC was introduced using LiDAR modality, methods based on camera images remain largely unexplored. Moreover, recent Transformer-based approaches utilize a fixed set of learned queries to reconstruct objects within the scene volume. Although these queries are typically updated with image context during training, they remain static at test time, limiting their ability to dynamically adapt specifically to the observed scene.
@@ 2 Related work
Semantic Scene Completion.   The task of Semantic Scene Completion has seen a large amount of interest after it was first introduced in lmscnet . The initial work focused on solving the task for indoor scenes from camera data zhang2018efficient ; zhang2019cascaded ; liu2018see ; li2019depth ; li2019rgbd ; li2020anisotropic ; chen20203d ; cai2021semantic and used point clouds in outdoor scenarios with LiDAR-based methods roldao20223d ; cheng2021s3cnet ; yan2021sparse ; li2021semisupervised ; rist2021semantic . In the following, we will discuss the most relevant contributions for Semantic Scene Completion in the context of autonomous driving. For a more thorough overview, we refer the reader to the surveys of xu2025survey and zhang2024vision .
@@
Recent progress in perception for autonomous driving has increasingly adopted Vision Transformer architectures dosovitskiyimage , notably DETR carion2020end . DETR introduced learnable object queries, which are trainable embeddings that allow the model to directly predict object sets in parallel, eliminating hand-crafted components like anchor boxes. In SSC, learnable queries can be adapted to predict semantic labels for 3D scene elements (e.g., voxels), streamlining the process and enhancing efficiency. Architectures like OccFormer occformer and VoxFormer li2023voxformer exemplify this trend, with OccFormer using multiscale voxel representations with query features to decode 3D semantic occupancy, and VoxFormer employing depth-based queries fused with image features, refined via deformable attention. Similarly, Symphonies symphonies refines instance queries iteratively using cross-attention between image features and 3D representations, while TPVFormer huang2023tpvformer leverages a tri-perspective view (TPV) to query scene features efficiently. HASSC wang2024not and CGFormer cgformer further refine this general approach by focusing on challenging voxels and contextual queries, respectively. Alternative approaches, such as implicit scenes via self-supervised learning in S4C, hayler2024s4c and CoHFF song2024cohff , which leverages both voxel and TPV representations to fuse features from multiple vehicles, have also emerged, though they do not utilize learnable queries.
@@
It is worth noting that existing literature presents ambiguity regarding context-awareness of queries for SSC. Although both Symphonies symphonies and CGFormer cgformer incorporate contextual cues during training, their approaches diverge at inference. In particular, Symphonies’ instance queries remain static and are shared across all input images, whereas CGFormer’s voxel queries preserve context-awareness by dynamically adapting to each input. To clarify this distinction, we differentiate between instance queries and instance proposals as context-aware and context-adaptive, respectively. Our introduced instance proposals effectively adapt scene characteristics to guide feature aggregation during both train and test time.
@@
Panoptic Scene Completion.   While panoptic LiDAR segmentation, which predicts panoptics for point clouds directly, has been widely studied Che2025S2MC-PHD ; Xiao2024lidarsegopenvoc ; xia2023lidarpanseg , voxel-based PSC approaches have only recently been introduced by PaSCo pasco . This method employs a hybrid convolutional neural network (CNN) and Transformer architecture with static instance queries to address LiDAR-based PSC, enhanced by an uncertainty-aware framework. In parallel, camera-based PSC from multi-frame and multi-view (surround-view) images is addressed by PanoOcc wang2024panoocc , which utilizes static voxel queries to aggregate spatiotemporal information, while SparseOcc liu2023sparseocc introduces a sparse voxel decoder and sparse static instance queries to predict semantic and instance occupancy from up to 96 temporal frames. Concurrently, PanoSSC shi2024panossc introduces a camera-based PSC method using forward-facing, non-temporal imagery. Its dual-head TPV-based architecture separates semantic occupancy and instance segmentation tasks to infer a 3D panoptic voxel grid. Likewise, our method targets PSC using forward-facing, non-temporal imagery. However, it overcomes a key limitation of prior approaches by avoiding reliance on static queries and instead introduces context-adaptive instance proposals that dynamically adapt to the scene at both train and test time. Notably, PanoSSC shi2024panossc is trained on a non-public post-processed version of the SemanticKITTI dataset semantickitti , which, in addition to the absence of released source code, challenges direct comparisons.
@@ 3.1 Overview
Given an input image $\mathbf{X}\in\mathbb{R}^{U\times V\times 3}$ (Fig. 2), the backbone predicts image features $\mathbf{F}\in\mathbb{R}^{H\times W\times F}$, and the depth net predicts a depth map $\mathbf{D}\in\mathbb{R}^{U\times V\times 1}$, following VoxFormer li2023voxformer . Further refining the image features and the depth map allows for generating context-aware 3D representations, denoted as 3D context features $\mathbf{F}_{\text{3D}}$ and initial voxels $\mathbf{V}$. Subsequently, visible and invisible voxels, sampled from $\mathbf{V}$ using the depth map, attend to the 3D image features $\mathbf{F}_{\text{3D}}$ via a series of 3D deformable cross and self-attention mechanisms. This process initializes instance proposals $\mathbf{I}_{P}$ and voxel proposals $\mathbf{V}_{P}$. In this setup, every instance proposal can eventually correlate with a single instance within the observed scene.
The voxel proposals $\mathbf{V}_{P}$ are encoded by 3D local and global encoding, in line with CGFormer cgformer , resulting in voxel features $\mathbf{V}_{F}$. We follow a two-stage training strategy in which the first stage uses these voxel features $\mathbf{V}_{F}$ to address SSC exclusively. This approach effectively guides the latent space towards semantics and geometry. The second training stage targets overall PSC, thereby detecting individual instances. This is achieved by attending the instance proposals $\mathbf{I}_{P}$ over the pre-trained voxel features from the first stage, resulting in instance features $\mathbf{I}_{F}$. These features are used to (i) predict semantics, and (ii) align instances and voxels to aggregate a complete panoptic scene.
@@
Following CGFormer cgformer , we refine the depth map $\mathbf{D}$ and the feature map $\mathbf{F}\in\mathbb{R}^{H\times W\times F}$ into per-pixel, frustum-shaped depth probabilities $\mathbf{D}_{R}\in\mathbb{R}^{H\times W\times D}$, effectively resulting in $|D|$ depth bins per pixel. This is achieved by applying $\Phi_{D}$ as a mixture of convolution and neighborhood cross-attention layers Hassani2023neigbor , formulated as $\mathbf{D}_{R}=\Phi_{D}(\mathbf{D},\mathbf{F}_{\text{2D}})$. Moreover, the generic image features $\mathbf{F}$ are projected into contextual features $\mathbf{F}_{\text{2D}}\in\mathbb{R}^{H\times W\times C}$ through the context net $\Phi_{C}$, denoted as $\mathbf{F}_{\text{2D}}=\Phi_{C}(\mathbf{F})$.
The next step is to lift context features $\mathbf{F}_{\text{2D}}$ to 3D, following the design of philion2020lift . To elaborate, we lift them to (i) 3D context features $\mathbf{F}_{\text{3D}}$ and (ii) 3D initial voxels $\mathbf{V}$. To achieve this, we distribute $\mathbf{F}_{\text{2D}}$ along rays defined by the camera intrinsics and weighted by the depth probability distribution $\mathbf{D}_{R}$. This is formulated as
@@
where $\mathcal{Q}$ quantizes continuous coordinates to discrete voxel indices, and $C$ represents the feature dimension. The summation aggregates all features that map to the same voxel, effectively accumulating evidence for stronger feature representations. Building upon the concept of cgformer , we add learnable embeddings to $\mathbf{V}$ that enhance its representational capacity.
@@ 3.3 Proposal Initialization
Voxel Visibility.   The initial voxels $\mathbf{V}$ are categorized into visible voxels $\mathbf{V}_{\text{vis}}$ and invisible voxels $\mathbf{V}_{\text{invis}}$, following established SSC works such as li2023voxformer . This is achieved by first applying inverse projection $\Pi^{-1}$ on the depth map $\mathbf{D}$ with camera intrinsics $\mathbf{K}\in\mathbb{R}^{4\times 4}$. This process generates a pseudo point cloud by unprojecting a pixel $(u,v)$ to a corresponding 3D point $(x,y,z)$. Furthermore, let $\mathbf{M}\in\mathbb{R}^{X\times Y\times Z}$ be a binary mask derived from this point cloud that filters the voxel features based on occupancy. Thus, the visible voxel features are defined as $\mathbf{V}_{\text{vis}}=\mathbf{V}\odot\mathbf{M}$, where $\mathbf{M}(x,y,z)=1$ if $(x,y,z)$ corresponds to an unprojected point, and 0 otherwise. The operator $\odot$ denotes the Schur product. Consequently, the invisible voxel features are given by $\mathbf{V}_{\text{invis}}=\mathbf{V}\odot(\mathbf{1}-\mathbf{M})$, with all-ones matrix $\mathbf{1}\in\mathbb{R}^{X\times Y\times Z}$. Note that this sampling does not include learnable parameters.
To further adapt $\mathbf{V}_{\text{vis}}$ with image context, we apply a tailored series of 3D deformable attention mechanisms, inspired by symphonies ; cgformer . Specifically, we apply 3D deformable cross-attention (DCA) to $\mathbf{V}_{\text{vis}}$ (as queries) and the 3D context features $\mathbf{F}_{\text{3D}}$ (as keys and values), resulting in updated visible voxel features $\tilde{\mathbf{V}}_{\text{vis}}^{\mathbf{x}}$. More precisely, an updated voxel query $\tilde{\mathbf{V}}_{\text{vis}}^{\mathbf{x}}$ at 3D location $\mathbf{x}$ computes
@@
Inspired by CGFormer´s cgformer query generator, our instance generator transforms the DSA output to initialize instance proposals via
@@
Voxel Proposals.   To initialize voxel proposals $\mathbf{V}_{P}$, we first merge the visible and invisible voxels through element-wise addition along dimension $C$, denoted as $\tilde{\mathbf{V}}_{\text{vis}}\oplus\mathbf{V}_{\text{invis}}$. We then apply 3D deformable self-attention to distribute updated context information over the entire scene volume, formulated as $\mathbf{V}_{P}=\text{DSA}(\tilde{\mathbf{V}}_{\text{vis}}\oplus\mathbf{V}_{\text{invis}})$, which is in line with li2023voxformer ; symphonies ; cgformer :

### v3

@@ 1 Introduction
Figure 1: Comparison of query initialization for Panoptic Scene Completion (PSC). Existing methods, e.g. Symphonies Jiang et al. (2024), randomly initialize instance queries and incorporate context-awareness during training. However, these queries retain their static nature at test time, as they are shared across all inputs. Our method IPFormer initializes them as instance proposals, which preserve context-adaptivity at test time, effectively aggregating directed features for improved PSC performance. Due to the the novelty of vision-based PSC and the absence of established baselines, we apply DBSCAN Ester et al. (1996) clustering to Symphonies’ SSC output to retrieve its individual instances.
@@
Panoptic Scene Completion provides a holistic scene understanding that can serve applications like autonomous driving and robotics by reconstructing volumetric geometry from sparse sensor data and assigning meaning to objects in the scene. Such holistic scene understanding requires both accurate geometry and semantic information about a scene. Historically, these tasks have been treated separately and evolved independently due to the distinct nature of geometric reconstruction and semantic interpretation, each requiring specialized algorithms and data representations. Despite the inherent interconnection between geometry and semantics, their separation hindered unified scene understanding. Classical structure-from-motion evolved into simultaneous localization and mapping (SLAM) systems for accurate reconstructions Mur-Artal and Tardos (2015); Engel et al. (2014); Koestler et al. (2022), while deep learning facilitated monocular depth estimation (MDE) methods Godard et al. (2017, 2019); Wimbauer et al. (2021), which are restricted to visible surfaces. In contrast, Scene Completion refers to the process of inferring the complete 3D structure of a scene, including both visible and occluded regions, from partial observations Wimbauer et al. (2023); Han et al. (2024).
@@
Apart from geometry, semantic image segmentation divides an image into semantically meaningful regions and labels them Yu and Koltun (2015); Chen et al. (2017). Panoptic segmentation Kirillov et al. (2019b, a); Porzi et al. (2019b) extends this approach by segmenting instances Hariharan et al. (2014) for things (countable objects like person or car) while maintaining semantic segmentation for stuff (amorphous regions like road or vegetation). This unified approach is vital for comprehensive scene understanding, as it distinguishes individual object instances while preserving semantic context for background elements, enabling applications like autonomous navigation. The recent trend in scene understanding has been to unify these tasks into Semantic Scene Completion (SSC) Hayler et al. (2024) and Panoptic Scene Completion (PSC) Cao et al. (2024). SSC integrates geometric completion with semantic labeling, predicting both the 3D structure and semantic categories, represented as a 3D grid composed of voxels. PSC builds on this by adding instance-level segmentation for each voxel, thus distinguishing between individual objects. While PSC was introduced using LiDAR modality, methods based on camera images remain largely unexplored. Moreover, recent Transformer-based approaches utilize a fixed set of learned queries to reconstruct objects within the scene volume. Although these queries are typically updated with image context during training, they remain static at test time, limiting their ability to dynamically adapt specifically to the observed scene.
@@ 2 Related work
Semantic Scene Completion.   The task of Semantic Scene Completion has seen a large amount of interest after it was first introduced in Roldão et al. (2020). The initial work focused on solving the task for indoor scenes from camera data Zhang et al. (2018, 2019); Liu et al. (2018); Li et al. (2019b, a, 2020); Chen et al. (2020); Cai et al. (2021) and used point clouds in outdoor scenarios with LiDAR-based methods Roldao et al. (2022); Cheng et al. (2021); Yan et al. (2021); Li et al. (2021); Rist et al. (2021). In the following, we will discuss the most relevant contributions for Semantic Scene Completion in the context of autonomous driving. For a more thorough overview, we refer the reader to the surveys of Xu et al. (2025) and Zhang et al. (2024).
@@
Recent progress in perception for autonomous driving has increasingly adopted Vision Transformer architectures Dosovitskiy et al. (2021), notably DETR Carion et al. (2020). DETR introduced learnable object queries, which are trainable embeddings that allow the model to directly predict object sets in parallel, eliminating hand-crafted components like anchor boxes. In SSC, learnable queries can be adapted to predict semantic labels for 3D scene elements (e.g., voxels), streamlining the process and enhancing efficiency. Architectures like OccFormer Zhang et al. (2023) and VoxFormer Li et al. (2023) exemplify this trend, with OccFormer using multiscale voxel representations with query features to decode 3D semantic occupancy, and VoxFormer employing depth-based queries fused with image features, refined via deformable attention. Similarly, Symphonies Jiang et al. (2024) refines instance queries iteratively using cross-attention between image features and 3D representations, while TPVFormer Huang et al. (2023) leverages a tri-perspective view (TPV) to query scene features efficiently. HASSC Wang et al. (2024a) and CGFormer Yu et al. (2024) further refine this general approach by focusing on challenging voxels and contextual queries, respectively. Alternative approaches, such as implicit scenes via self-supervised learning in S4C, Hayler et al. (2024) and CoHFF Song et al. (2024), which leverages both voxel and TPV representations to fuse features from multiple vehicles, have also emerged, though they do not utilize learnable queries.
@@
It is worth noting that existing literature presents ambiguity regarding context-awareness of queries for SSC. Although both Symphonies Jiang et al. (2024) and CGFormer Yu et al. (2024) incorporate contextual cues during training, their approaches diverge at inference. In particular, Symphonies’ instance queries remain static and are shared across all input images, whereas CGFormer’s voxel queries preserve context-awareness by dynamically adapting to each input. To clarify this distinction, we differentiate between instance queries and instance proposals as context-aware and context-adaptive, respectively. Our introduced instance proposals effectively adapt scene characteristics to guide feature aggregation during both train and test time.
@@
Panoptic Scene Completion.   While panoptic LiDAR segmentation, which predicts panoptics for point clouds directly, has been widely studied Chen et al. (2025); Xiao et al. (2025); Xiang et al. (2023), voxel-based PSC approaches have only recently been introduced by PaSCo Cao et al. (2024). This method employs a hybrid convolutional neural network (CNN) and Transformer architecture with static instance queries to address LiDAR-based PSC, enhanced by an uncertainty-aware framework. In parallel, camera-based PSC from multi-frame and multi-view (surround-view) images is addressed by PanoOcc Wang et al. (2024b), which utilizes static voxel queries to aggregate spatiotemporal information, while SparseOcc Liu et al. (2023) introduces a sparse voxel decoder and sparse static instance queries to predict semantic and instance occupancy from up to 96 temporal frames. Concurrently, PanoSSC Shi et al. (2024) introduces a camera-based PSC method using forward-facing, non-temporal imagery. Its dual-head TPV-based architecture separates semantic occupancy and instance segmentation tasks to infer a 3D panoptic voxel grid. Likewise, our method targets PSC using forward-facing, non-temporal imagery. However, it overcomes a key limitation of prior approaches by avoiding reliance on static queries and instead introduces context-adaptive instance proposals that dynamically adapt to the scene at both train and test time. Notably, PanoSSC Shi et al. (2024) is trained on a non-public post-processed version of the SemanticKITTI dataset Behley et al. (2019), which, in addition to the absence of released source code, challenges direct comparisons.
@@ 3.1 Overview
Given an input image $\mathbf{X}\in\mathbb{R}^{U\times V\times 3}$ (Fig. 2), the backbone predicts image features $\mathbf{F}\in\mathbb{R}^{H\times W\times F}$, and the depth net predicts a depth map $\mathbf{D}\in\mathbb{R}^{U\times V\times 1}$, following VoxFormer Li et al. (2023). Further refining the image features and the depth map allows for generating context-aware 3D representations, denoted as 3D context features $\mathbf{F}_{\text{3D}}$ and initial voxels $\mathbf{V}$. Subsequently, visible and invisible voxels, sampled from $\mathbf{V}$ using the depth map, attend to the 3D image features $\mathbf{F}_{\text{3D}}$ via a series of 3D deformable cross and self-attention mechanisms. This process initializes instance proposals $\mathbf{I}_{P}$ and voxel proposals $\mathbf{V}_{P}$. In this setup, every instance proposal can eventually correlate with a single instance within the observed scene.
The voxel proposals $\mathbf{V}_{P}$ are encoded by 3D local and global encoding, in line with CGFormer Yu et al. (2024), resulting in voxel features $\mathbf{V}_{F}$. We follow a two-stage training strategy in which the first stage uses these voxel features $\mathbf{V}_{F}$ to address SSC exclusively. This approach effectively guides the latent space towards semantics and geometry. The second training stage targets overall PSC, thereby detecting individual instances. This is achieved by attending the instance proposals $\mathbf{I}_{P}$ over the pre-trained voxel features from the first stage, resulting in instance features $\mathbf{I}_{F}$. These features are used to (i) predict semantics, and (ii) align instances and voxels to aggregate a complete panoptic scene.
@@
Following CGFormer Yu et al. (2024), we refine the depth map $\mathbf{D}$ and the feature map $\mathbf{F}\in\mathbb{R}^{H\times W\times F}$ into per-pixel, frustum-shaped depth probabilities $\mathbf{D}_{R}\in\mathbb{R}^{H\times W\times D}$, effectively resulting in $|D|$ depth bins per pixel. This is achieved by applying $\Phi_{D}$ as a mixture of convolution and neighborhood cross-attention layers Hassani et al. (2023), formulated as $\mathbf{D}_{R}=\Phi_{D}(\mathbf{D},\mathbf{F}_{\text{2D}})$. Moreover, the generic image features $\mathbf{F}$ are projected into contextual features $\mathbf{F}_{\text{2D}}\in\mathbb{R}^{H\times W\times C}$ through the context net $\Phi_{C}$, denoted as $\mathbf{F}_{\text{2D}}=\Phi_{C}(\mathbf{F})$.
The next step is to lift context features $\mathbf{F}_{\text{2D}}$ to 3D, following the design of Philion and Fidler (2020). To elaborate, we lift them to (i) 3D context features $\mathbf{F}_{\text{3D}}$ and (ii) 3D initial voxels $\mathbf{V}$. To achieve this, we distribute $\mathbf{F}_{\text{2D}}$ along rays defined by the camera intrinsics and weighted by the depth probability distribution $\mathbf{D}_{R}$. This is formulated as
@@
where $\mathcal{Q}$ quantizes continuous coordinates to discrete voxel indices, and $C$ represents the feature dimension. The summation aggregates all features that map to the same voxel, effectively accumulating evidence for stronger feature representations. Building upon the concept of Yu et al. (2024), we add learnable embeddings to $\mathbf{V}$ that enhance its representational capacity.
@@ 3.3 Proposal Initialization
Voxel Visibility.   The initial voxels $\mathbf{V}$ are categorized into visible voxels $\mathbf{V}_{\text{vis}}$ and invisible voxels $\mathbf{V}_{\text{invis}}$, following established SSC works such as Li et al. (2023). This is achieved by first applying inverse projection $\Pi^{-1}$ on the depth map $\mathbf{D}$ with camera intrinsics $\mathbf{K}\in\mathbb{R}^{4\times 4}$. This process generates a pseudo point cloud by unprojecting a pixel $(u,v)$ to a corresponding 3D point $(x,y,z)$. Furthermore, let $\mathbf{M}\in\mathbb{R}^{X\times Y\times Z}$ be a binary mask derived from this point cloud that filters the voxel features based on occupancy. Thus, the visible voxel features are defined as $\mathbf{V}_{\text{vis}}=\mathbf{V}\odot\mathbf{M}$, where $\mathbf{M}(x,y,z)=1$ if $(x,y,z)$ corresponds to an unprojected point, and 0 otherwise. The operator $\odot$ denotes the Schur product. Consequently, the invisible voxel features are given by $\mathbf{V}_{\text{invis}}=\mathbf{V}\odot(\mathbf{1}-\mathbf{M})$, with all-ones matrix $\mathbf{1}\in\mathbb{R}^{X\times Y\times Z}$. Note that this sampling does not include learnable parameters.
To further adapt $\mathbf{V}_{\text{vis}}$ with image context, we apply a tailored series of 3D deformable attention mechanisms, inspired by Jiang et al. (2024); Yu et al. (2024). Specifically, we apply 3D deformable cross-attention (DCA) to $\mathbf{V}_{\text{vis}}$ (as queries) and the 3D context features $\mathbf{F}_{\text{3D}}$ (as keys and values), resulting in updated visible voxel features $\tilde{\mathbf{V}}_{\text{vis}}^{\mathbf{x}}$. More precisely, an updated voxel query $\tilde{\mathbf{V}}_{\text{vis}}^{\mathbf{x}}$ at 3D location $\mathbf{x}$ computes
@@
Inspired by CGFormer´s Yu et al. (2024) query generator, our instance generator transforms the DSA output to initialize instance proposals via
@@
Voxel Proposals.   To initialize voxel proposals $\mathbf{V}_{P}$, we first merge the visible and invisible voxels through element-wise addition along dimension $C$, denoted as $\tilde{\mathbf{V}}_{\text{vis}}\oplus\mathbf{V}_{\text{invis}}$. We then apply 3D deformable self-attention to distribute updated context information over the entire scene volume, formulated as $\mathbf{V}_{P}=\text{DSA}(\tilde{\mathbf{V}}_{\text{vis}}\oplus\mathbf{V}_{\text{invis}})$, which is in line with Li et al. (2023); Jiang et al. (2024); Yu et al. (2024):